Riza is 25 years old while her brother was born. After the sum of their ages is 105. What is the age of riza and her brother today?
Answer by ankor@dixie-net.com(22740)

let r = R's age
let b = bro's age
:
Riza is 25 years old while her brother was born.
From this statement we can say:
r = b + 25
:
Now the sum of their ages is 105, what is their present age?
r + b = 105
replace r with (b+25), find b
(b+25) + b = 105
2b = 105 - 25
2b = 80
b = 80/2
b = 40 yr is brother's present age
then
25 + 40 = 65 yrs is Riza's present age
:
:
Check this by adding up their ages, and seeing it equals 105